what volume of a 25 muM solution of MgCl would contain 5 pmol of MgCl?
Answer #1

pmol = pico mol = 10^(-12) mol

25 muM = 25 * 10^(-6) M solution

1 L of MgCl solution will contain 25 * 10^(-6) moles of MgCl

Volume of solution(required for 5 pmol) = [5 * 10^(-12)]/[25 * 10^(-6)] * 1 L = 2 * 10^(-7) L or 0.2 uL

Hence we need to use 0.2 uL of 25 muM solution of MgCl for getting 5 pico mol of MgCl

Note - Post any doubts/queries in comments section.